OpenStudy (anonymous):

Help??? Simplify and state restrictions: 12x^2-25x+12/3x^2+2X-8

OpenStudy (anonymous):

factorise the numerator and denominator and cancel the common factor.

OpenStudy (anonymous):

I'm hving trouble coming up with the factored numerator, though. @surjithayer

OpenStudy (anonymous):

3*-8=-24 6-4=2 6*-4=-24 write 2x=6x-4x

OpenStudy (anonymous):

for numerator 12*12=144 -16-9=-25 -16*-9=144 write-25x=-16x-9x

OpenStudy (anonymous):

So, for the numerator is would be (12x-16)(x-9)?? @surjithayer

OpenStudy (jdoe0001):